OpenStudy (anonymous):

After the fraction x plus 1 all over 2 minus the fraction x plus 2 all over 3 x have been combined using the Least Common Denominator of 6x, what is the numerator?

OpenStudy (doc.brown):

\[\frac{x+1}{2}-\frac{x+2}{3x}\]\[\frac{3x}{3x}(\frac{x+1}{2})-\frac{2}{2}(\frac{x+2}{3x})\]\[\frac{3x^{2}+3x}{6x}-\frac{2x+4}{6x}\]\[\frac{3x^{2}+3x-2x-4}{6x}\] And the numerator is\[3x^{2}+x-4\]